Meridional number of a link and shrinkability of toroidal decompositions

Jože Malešič, Ljubljana, Slovenia

Keywords: link, solid torus, toroidal decomposition, shrinkability, Whitehead–Bing decomposition, satellite of link, meridional number AMS Subject classification: 57M25, 54B15

Abstract: An invariant for embeddings of links into the solid torus, called the meridional number, is introduced. Techniques are developed which produce a lower estimate of the meridional number in the case when the given link can be expressed as a satellite of another link. Such estimates yield a sufficient condition for shrinkability of toroidal decomposition.